# Statistical Inference of Constrained Stochastic Optimization via Sketched Sequential Quadratic Programming

### v4

@@ 3 Adaptive Inexact StoSQP Method
We combine the above three steps and summarize AI-StoSQP in Algorithm 1. To end this section, we introduce a filtration notation for later use. We let $\mathcal{F}_{t}=\sigma(\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t})$, $\forall t\geq 0$ be the $\sigma$-algebra generated by the random variables $\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t}$. Moreover, we let $\mathcal{F}_{t-2/3}=\sigma(\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t-1}\cup\xi_{t})$, $\mathcal{F}_{t-1/3}=\sigma(\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t-1}\cup\xi_{t}\cup\{S_{t,j}\}_{j})$, and have $\mathcal{F}_{t-1}\subseteq\mathcal{F}_{t-2/3}\subseteq\mathcal{F}_{t-1/3}\subseteq\mathcal{F}_{t}$. For consistency, $\mathcal{F}_{-1}$ is the trivial $\sigma$-algebra. With these notation, Algorithm 1 has a generating process as follows: given $({\bm{x}}_{t},{\bm{\lambda}}_{t})$, we first realize $\xi_{t}$ to estimate the gradient ${\bar{g}}_{t}$ and Hessian ${\bar{H}}_{t}$ and derive $\mathcal{F}_{t-2/3}$; then we generate $\{S_{t,j}\}_{j}$ to obtain the inexact Newton direction and derive $\mathcal{F}_{t-1/3}$; then we select the stepsize ${\bar{\alpha}}_{t}$ and derive $\mathcal{F}_{t}$. We also let $(\Delta{\bm{x}}_{t},\Delta{\bm{\lambda}}_{t})$ be the exact Newton direction solved from (3.2) with ¯${\bar{\nabla}}_{{\bm{x}}}\mathcal{L}_{t}$ being replaced by $\nabla_{{\bm{x}}}\mathcal{L}_{t}$.
@@ 4 Global Almost Sure Convergence
In this section, we present the global almost sure convergence [^1] Global convergence in nonlinear optimization refers to the convergence to a stationary point from any initialization, in contrast to the convergence to a global solution, which is not achievable without particular problem structures (Nocedal2006Numerical). However, they are equivalent for convex problems as studied for projection-based methods in Duchi2021Asymptotic; Davis2023Asymptotic. for the StoSQP method. We show that the KKT residual $\|\nabla\mathcal{L}_{t}\|$ converges to zero from any initialization. This global convergence serves as a preliminary result of our inference analysis in Section LABEL:sec:4.
@@
$\mathcal{L}_{\mu,\nu}({\bm{x}},{\bm{\lambda}})=\mathcal{L}({\bm{x}},{\bm{\lambda}})+\frac{\mu}{2}\|c({\bm{x}})\|^{2}+\frac{\nu}{2}\|\nabla_{{\bm{x}}}\mathcal{L}({\bm{x}},{\bm{\lambda}})\|^{2},\quad\quad\text{with }\mu,\nu>0.$ (4.1)
The first penalty term biases the feasibility error; while, in contrast to standard augmented Lagrangian ($\nu=0$), the extra second penalty term biases the optimality error. We will first show that the inner product between the exact Newton direction $(\Delta{\bm{x}}_{t},\Delta{\bm{\lambda}}_{t})$ and the augmented Lagrangian gradient $\nabla\mathcal{L}_{\mu,\nu}$, with proper parameters $\mu$ and $\nu$, is sufficiently negative (Lemma LABEL:lem:2). Thus, the Newton direction is a descent direction of $\mathcal{L}_{\mu,\nu}$. Then, we will show the augmented Lagrangian $\mathcal{L}_{\mu,\nu}$ decreases at each step even with an inexact Newton direction (Lemma LABEL:lem:3), so that the residual $\|\nabla\mathcal{L}_{t}\|$ finally vanishes to zero (Theorem LABEL:thm:1).
@@ 4.1 Assumptions and preliminary results
Assumption 4.1 assumes $G_{t}$ has full row rank, which is referred to as the linear independence constraint qualification (LICQ). LICQ is a common constraint qualification ensuring the uniqueness of the dual solution, and is also necessary for the inference analysis (see (Duchi2021Asymptotic, Assumption B) and (Davis2023Asymptotic, Example 2.1)). LICQ and conditions on $B_{t}$ are critical for SQP methods; they imply QP (3.1) has a unique solution (Nocedal2006Numerical, Lemma 16.1). The Lipschitz continuity of the Hessian matrix is also standard for analyzing Newton’s method.

### v5

@@ 3 Adaptive Inexact StoSQP Method
We combine the above three steps and summarize AI-StoSQP in Algorithm 1. To end this section, we introduce a filtration notation for later use. We let $\mathcal{F}_{t}=\sigma(\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t})$, $\forall t\geq 0$ be the $\sigma$-algebra generated by the random variables $\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t}$. Moreover, we let $\mathcal{F}_{t-2/3}=\sigma(\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t-1}\cup\xi_{t})$, $\mathcal{F}_{t-1/3}=\sigma(\{\xi_{i},\{S_{i,j}\}_{j},{\bar{\alpha}}_{i}\}_{i=0}^{t-1}\cup\xi_{t}\cup\{S_{t,j}\}_{j})$, and have $\mathcal{F}_{t-1}\subseteq\mathcal{F}_{t-2/3}\subseteq\mathcal{F}_{t-1/3}\subseteq\mathcal{F}_{t}$. For consistency, $\mathcal{F}_{-1}$ is the trivial $\sigma$-algebra. With these notation, Algorithm 1 has a generating process as follows: given $({\bm{x}}_{t},{\bm{\lambda}}_{t})$, we first realize $\xi_{t}$ to get the estimates of the gradient ${\bar{g}}_{t}$ and Hessian ${\bar{H}}_{t}$ and derive $\mathcal{F}_{t-2/3}$; then we generate $\{S_{t,j}\}_{j}$ to obtain the inexact Newton direction and derive $\mathcal{F}_{t-1/3}$; then we select the stepsize ${\bar{\alpha}}_{t}$ and derive $\mathcal{F}_{t}$. We also let $(\Delta{\bm{x}}_{t},\Delta{\bm{\lambda}}_{t})$ be the exact Newton direction solved from (3.2) with ¯${\bar{\nabla}}_{{\bm{x}}}\mathcal{L}_{t}$ being replaced by $\nabla_{{\bm{x}}}\mathcal{L}_{t}$.
@@ 4 Global Almost Sure Convergence
In this section, we present the global almost sure convergence [^1] Global convergence in nonlinear optimization refers to the convergence to a stationary point from any initialization, in contrast to the convergence to a global solution, which is not achievable without particular problem structures (Nocedal2006Numerical). However, they are equivalent for convex problems as studied for projection-based methods in Duchi2021Asymptotic; Davis2024Asymptotic. for the StoSQP method. We show that the KKT residual $\|\nabla\mathcal{L}_{t}\|$ converges to zero from any initialization. This global convergence serves as a preliminary result of our inference analysis in Section LABEL:sec:4.
@@
$\mathcal{L}_{\mu,\nu}({\bm{x}},{\bm{\lambda}})=\mathcal{L}({\bm{x}},{\bm{\lambda}})+\frac{\mu}{2}\|c({\bm{x}})\|^{2}+\frac{\nu}{2}\|\nabla_{{\bm{x}}}\mathcal{L}({\bm{x}},{\bm{\lambda}})\|^{2},\quad\quad\text{with }\mu,\nu>0.$
The first penalty term biases the feasibility error; while, in contrast to standard augmented Lagrangian ($\nu=0$), the extra second penalty term biases the optimality error. We will first show that the inner product between the exact Newton direction $(\Delta{\bm{x}}_{t},\Delta{\bm{\lambda}}_{t})$ and the augmented Lagrangian gradient $\nabla\mathcal{L}_{\mu,\nu}$, with proper parameters $\mu$ and $\nu$, is sufficiently negative (Lemma LABEL:lem:2). Thus, the Newton direction is a descent direction of $\mathcal{L}_{\mu,\nu}$. Then, we will show that the augmented Lagrangian $\mathcal{L}_{\mu,\nu}$ decreases at each step even with an inexact Newton direction (Lemma LABEL:lem:3). This implies that the residual $\|\nabla\mathcal{L}_{t}\|$ finally vanishes to zero (Theorem LABEL:thm:1).
@@ 4.1 Assumptions and preliminary results
Assumption 4.1 assumes $G_{t}$ has full row rank, which is referred to as the linear independence constraint qualification (LICQ). LICQ is a common constraint qualification ensuring the uniqueness of the dual solution, and is also necessary for the inference analysis (see (Duchi2021Asymptotic, Assumption B) and (Davis2024Asymptotic, Example 2.1)). LICQ and conditions on $B_{t}$ are critical for SQP methods; they imply QP (3.1) has a unique solution (Nocedal2006Numerical, Lemma 16.1). The Lipschitz continuity of the Hessian matrix is also standard for analyzing Newton’s method.
The bounded iterates condition is commonly assumed in the literature on (stochastic) nonlinear nonconvex optimization, both for first-order gradient-based methods (Bolte2013Proximal; Song2014Weak; Davis2016sound; Davis2019Stochastic; Atchade2017perturbed; Asi2019Stochastic; Liu2023Statistical) and for second-order SQP methods ((Bertsekas1982Constrained, Proposition 4.15), (Nocedal2006Numerical, Theorem 18.3)). This assumption ensures that all functions with their gradients and Hessians are bounded over $\mathcal{X}\times\Lambda$ as long as they are smooth. Some literature replaces the bounded iterates condition by directly imposing boundedness on the gradients and Hessians of the objective and constraints, although the main use of the condition in the proof is rather similar (Berahas2021Sequential; Berahas2023Stochastic; Curtis2021Inexact; Ramprasad2022Online; Liu2023Online).
We provide two justifications for the boundedness condition. First, in our study, the StoSQP iterates presumably track a deterministic feasible set $\{{\bm{x}}\in\mathbb{R}^{d}:c({\bm{x}})={\bm{0}}\}$, so we believe that an unbounded iteration sequence is generally rare especially when the feasible set is bounded. Second, a practical way to enforce the boundedness condition may be through adaptive truncation (Andrieu2005Stability; Liang2010Trajectory). Under some conditions on the Markov transition kernel of the iteration sequence, one can show that the truncation occurs only finitely many times, ensuring that the convergence and asymptotic behavior are finally not affected by the truncation.